# Collectively enhanced ground-state cooling in subwavelength atomic arrays

### v2

@@ E.2 Dependence on the axis of atomic motion
In the main text, we demonstrate enhanced cooling of interacting chains of atoms located on the $x$-axis and moving along the perpendicular direction, that is, along the $z$-axis. Then, the first-order spin-motion interaction in $\eta$, given by Eq. (59a), receives contributions solely from the laser drive, and is proportional to $\Omega_{j}^{\prime}$. In this section, we extend the analysis to atomic motion along the chain, i. e., along the $x$-axis, such that the spin-motion coupling acquires an additional contribution from dipole-dipole interactions, proportional to $G_{ij}^{\prime}$.
@@
and similarly for $s_{2}^{(0)}$. For $\bar{\nu}\gg|J_{12}|,|J_{12}^{\prime}|,\gamma_{12},\gamma_{12}^{\prime}$, the displacement can be approximated to $s_{j}^{(0)}\approx-\Omega_{j}/\bar{\nu}$. Then, the strength of the spin-motion coupling arising from light-induced dipole-dipole interactions is much smaller than the contribution from the driving field, i. e., $\eta_{j}|J_{12}^{\prime}||s_{j}^{(0)}|\approx\eta_{j}|\Omega_{j}||J_{12}^{\prime}|/\bar{\nu}\ll\eta_{j}|\Omega_{j}|$, and the steady state phonon number $\bar{n}$ is identical for parallel and perpendicular motion. This is shown in Fig. 10(c), where we compare $\bar{n}$ for parallel motion obtained by solving the full master equation (7) including the spin and motional degrees of freedom (markers) with the $\bar{n}$ for perpendicular motion obtained by solving the effective master equation (1). For $|J_{12}^{\prime}|\gtrsim\bar{\nu}$, the dipole-dipole contribution to the spin-motion interaction can become dominant, and the steady state phonon number for parallel and out-of-plane motion can differ. As shown in Fig. 10(c), this divergence occurs at larger $\bar{\nu}$ for smaller atomic separations $d$ as $|J_{12}^{\prime}|$ increases. Notably, enhanced cooling is still found for a wide range of trap frequencies $\bar{\nu}$ and distances $d$.
Figure 10: (a) Two trapped atoms at a distance $d$ along the $x$-axis move in the perpendicular direction ($z$-axis). The drive is applied along the direction of motion, such that $\Omega_{j}=\Omega e^{ik_{0}z_{j}}\equiv\Omega$. (b) Two trapped atoms at a distance $d$ along the $x$-axis move in the same direction (parallel motion). The drive is applied along the direction of motion, such that $\Omega_{j}=\Omega e^{ik_{0}x_{j}}$. (c) Steady state phonon number per atom $\bar{n}$ for perpendicular (dashed lines) and parallel (markers) motion, respectively obtained by solving the effective master equation (1) and the full master equation (7) including spin and motion. We consider $\eta=0.02$ and $\delta\nu=10^{-3}\gamma_{0}$.
@@ E.3 Coupling to other directions of motion
We consider the same setup as in the main text: The atoms are arranged along the $x$-axis and driven by a classical plane-wave field propagating along $\mathbf{z}$, with drive parameters $\Omega_{j}=\Omega$ and $\Omega_{j}^{(\alpha)}=i\Omega\delta_{\alpha z}$. In this configuration the laser directly cools only the motion along the $z$-axis. The dissipative and coherent couplings between motion along different directions arise from two different processes. The first is recoil heating, and vanishes since $G_{ij}^{\alpha\beta}=0$ for all pairs of directions $\alpha\neq\beta$ [see Eq. (C.2)]. The second is the interference between laser- and dipole-induced spin-motion coupling to first-order in $\eta$, given by Eq. (C.1). Notably, $\Omega_{j}^{(\alpha)}$ and $G_{ij}^{(\beta)}$ are only non-zero for $\alpha=z$ and $\beta=x$, respectively. Thus, motion along $\mathbf{z}$ only couples to motion along $\mathbf{x}$ (through terms of the form $\Omega_{j}^{(z)}G_{ij}^{(x)}$), while motion along $\mathbf{y}$ remains decoupled. For trap frequencies $\nu\gg\gamma_{0}$, this residual coupling is much smaller than the cooling rates along $\mathbf{z}$, $\mathcal{R}_{iziz}^{-}$, and thus it does not impact the cooling dynamics. When $\nu\lesssim\gamma_{0}$, such couplings can still be suppressed under the rotating wave approximation by applying slightly different trap frequencies along $\mathbf{x}$ and $\mathbf{z}$. In this regime, motional degrees of freedom along different directions evolve independently, and enhanced cooling along $\mathbf{z}$ remains possible under the array cooling conditions discussed in the main text. On the other hand, motion along $\mathbf{x}$ and $\mathbf{y}$ can still experience heating, either from recoil (for both directions) or from dipole-induced spin-motion coupling (for $\mathbf{x}$). However, since these heating rates are typically much smaller than the cooling rate along $\mathbf{z}$, they do not lead to appreciable heating during the cooling process.
Figure 11: Two atoms separated by $d=0.2\lambda_{0}$ along the $x$-axis, polarized along the $y$-axis and subject to a three-dimensional harmonic confinement with trapping frequencies $\nu_{1x}=\nu_{1y}=\nu_{1z}=20\gamma_{0}$ for the first atom and $\nu_{1x}=\nu_{1y}=\nu_{1z}=20.01\gamma_{0}$ for the second atom. A drive with Rabi frequency $\Omega=0.1\gamma_{0}$ is applied along the $z$ direction, and we further consider $\eta=0.02$. We plot the average photon number over time along all three directions of motion, $\bar{n}_{\mathrm{\alpha,3D}}$ with $\alpha\in\{x,y,z\}$. The phonon number along the $z$ direction, $\bar{n}_{\mathrm{z,3D}}$, is identical to the case where motion along the $x$ and $y$ directions are neglected, $\bar{n}_{\mathrm{z,1D}}$.
@@ Appendix F Numerical optimization of the cooling rate
Figure 12: (a) Steady state phonon number $\bar{n}$ and (b) cooling rate $\Gamma$ as a function of drive strength $\Omega$ for two atoms separated by a distance $d=0.2\lambda_{0}$ along the $x$-axis, polarized along the $y$-axis and moving along the $z$-axis. The inset shows an instance of the exponentially decaying phonon number over time for $\Omega=0.2\gamma_{0}$, from which the cooling rate is fitted. (c) Steady state phonon number versus cooling rate for various spacings. Each point corresponds to a different value of $\Omega$. The squared marker corresponds to the critical cooling rate $\Gamma_{c}$, obtained at the drive strength that results in a steady state phonon number $10\%$ larger than the optimal one. The triangular marker corresponds to the alternative critical cooling rate $\Gamma_{c}^{\prime}$, obtained at the drive that results in a largest cooling rate while mantaining a steady state phonon number smaller than $1.1\bar{n}_{\textit{ind}}$. (d-e) Critical cooling rates $\Gamma_{c}$ and $\Gamma_{c}^{\prime}$ and their corresponding steady state phonon numbers for chains of $N$ atoms and various spacings. In all panels, we consider $\eta=0.02$ and $\nu_{j}=\bar{\nu}+\delta\nu(j-\lfloor N/2\rfloor-1)$ with $\bar{\nu}=20\gamma_{0}$ and $\delta\nu=5\times 10^{-3}\gamma_{0}$.
@@
Let us consider as an example two atoms in the array cooling regime as defined in Eq. (32) and where the subradiant cooling sideband are resolved (i. e., $\gamma_{A}\ll\bar{\nu}$). The steady state phonon number $\bar{n}$ and the cooling rate $\Gamma$ at which it is reached depend on the drive strength. For sufficiently small drive strength, such that the system remains in the array cooling regime $\Omega^{2}\ll\delta\nu\gamma_{A}/2\eta^{2}$ and the spin-motion coupling is weak $\Omega\ll\gamma_{A}/\eta$, the phonon number exponentially decays to its steady state value at a rate $\Gamma=\mathcal{R}^{-}-\mathcal{R}^{+}\approx\mathcal{R}^{-}\approx 2\eta^{2}\Omega^{2}/\gamma_{A}$, as shown in Fig. 12(a) and (b) for $\bar{\nu}=20\gamma_{0}$. As the drive strength increases, either the array cooling condition or the weak spin-motion coupling condition breaks down, and the effective master equation resulting in Eq. (16) stops being valid. In that case, $\bar{n}$ and $\Gamma$ can only be obtained by numerically solving the full system including spin and motional degrees of freedom. More precisely, $\Gamma$ is obtained by fitting an exponential decay to the phonon number over time, as shown in the inset of Fig. 12(a). We find that, as soon as one of the two conditions is violated, the phonon number starts to increase with increasing $\Omega$ and the cooling rate eventually decreases. To better understand this process, we plot $\bar{n}$ as a function of cooling rate in Fig. 12(c). That is, each data point in this plot corresponds to a different value of $\Omega$. In the main text, we defined the critical cooling rate $\Gamma_{c}$ as the one obtained at the driving strength $\Omega$ that leads to a steady state phonon number $10\%$ larger than the optimal one. For two atoms at a distance $d=0.1\lambda_{0}$, this corresponds to the squared marker in Fig. 12(c). However, this condition is more restrictive for lattices with darker spin modes than attain lower phonon numbers. In other words, we could drive the two atoms at a distance $d=0.1\lambda_{0}$ stronger and thereby attain a larger cooling rate while still reaching steady state phonon numbers smaller than in the case of independent atoms, $\bar{n}<\bar{n}_{\textit{ind}}$. This motivates us to introduce an alternative definition for the critical cooling rate $\Gamma_{c}^{\prime}$ as the largest possible rate that can be reached such that the phonon number remains smaller than $1.1\bar{n}_{\textit{ind}}$. As an example, $\Gamma_{c}^{\prime}$ and its corresponding $\bar{n}^{\prime}$ are shown by the triangular marker in Fig. 12(c) for $d=0.1\lambda_{0}$. In Fig. 12(d) and (e), we plot $\bar{n}$ and $\Gamma_{c}$ for chains of $N$ atoms with various spacing $d$ for $\bar{\nu}=20\gamma_{0}$. Again, squared and triangular markers denote the two different definitions for the critical cooling rate. In all cases, $\Gamma_{c}^{\prime}\gtrsim 2\Gamma_{c}$, showing that the critical cooling rates presented in the main text can be sped up by driving the system stronger at the expense of reaching larger steady state phonon numbers $\bar{n}^{\prime}$. This suggests the following protocol for efficient cooling: first, the system is cooled at $\Gamma^{\prime}$ to a phonon number close to $\bar{n}^{\prime}$; then, the drive is dynamically reduced until the lowest possible phonon number $\bar{n}$ is eventually reached.
@@ Appendix G Cooling $N$ atoms in different geometries
Figure 13: Steady state phononon number $\bar{n}$ for arrays of different geometries as a function of atom number $N$, for (a) $\bar{\nu}=20\gamma_{0}$ and (b) $\bar{\nu}=0.25\gamma_{0}$. We consider a linear ordered chain, a square two-dimensional lattice and a ring of atoms. The atomic chain is placed along the $x$-axis and the atoms are polarized along the $y$-direction. In all cases, the atoms move along the $z$-direction. The squared array and the ring are placed in the $xy$-plane and the atoms are circularly polarized in the $xy$-plane. For the atomic chain and the squared array, we consider a lattice spacing $d=0.2\lambda_{0}$, while the ring has a radius such that neighboring atoms are at a distance $d=0.2\lambda_{0}$. The results are obtained for $\eta=0.02$ and all atoms have different trap frequencies in increments of $\delta\nu=5\times 10^{-4}\gamma_{0}$, such that the system is in the array cooling regime.

### v3

@@ E.2 Dependence on the axis of atomic motion
In the main text, we demonstrate enhanced cooling of interacting chains of atoms located on the $x$ axis and moving along the perpendicular direction, that is, along the $z$ axis. Then, the first-order spin-motion interaction in $\eta$, given by Eq. (59a), receives contributions solely from the laser drive, and is proportional to $\Omega_{j}^{\prime}$. In this section, we extend the analysis to atomic motion along the chain, i. e., along the $x$ axis, such that the spin-motion coupling acquires an additional contribution from dipole-dipole interactions, proportional to $G_{ij}^{\prime}$.
@@
and similarly for $s_{2}^{(0)}$. For $\bar{\nu}\gg|J_{12}|,|J_{12}^{\prime}|,\gamma_{12},\gamma_{12}^{\prime}$, the displacement can be approximated to $s_{j}^{(0)}\approx-\Omega_{j}/\bar{\nu}$. Then, the strength of the spin-motion coupling arising from light-induced dipole-dipole interactions is much smaller than the contribution from the driving field, i. e., $\eta_{j}|J_{12}^{\prime}||s_{j}^{(0)}|\approx\eta_{j}|\Omega_{j}||J_{12}^{\prime}|/\bar{\nu}\ll\eta_{j}|\Omega_{j}|$, and the steady state phonon number $\bar{n}$ is identical for parallel and perpendicular motion. This is shown in Fig. 10(c), where we compare $\bar{n}$ for parallel motion obtained by solving the full master equation (7) including the spin and motional degrees of freedom (markers) with $\bar{n}$ for perpendicular motion obtained by solving the effective master equation (1). For $|J_{12}^{\prime}|\gtrsim\bar{\nu}$, the dipole-dipole contribution to the spin-motion interaction can become dominant, and the steady state phonon number for parallel and out-of-plane motion can differ. As shown in Fig. 10(c), this divergence occurs at larger $\bar{\nu}$ for smaller atomic separations $d$ as $|J_{12}^{\prime}|$ increases. Notably, enhanced cooling is still found for a wide range of trap frequencies $\bar{\nu}$ and distances $d$.
Figure 10: (a) Two trapped atoms at a distance $d$ along the $x$ axis move in the perpendicular direction ($z$ axis). The drive is applied along the direction of motion, such that $\Omega_{j}=\Omega e^{ik_{0}z_{j}}\equiv\Omega$. (b) Two trapped atoms at a distance $d$ along the $x$-axis move in the same direction (parallel motion). The drive is applied along the direction of motion, such that $\Omega_{j}=\Omega e^{ik_{0}x_{j}}$. (c) Steady-state phonon number per atom $\bar{n}$ for perpendicular (dashed lines) and parallel (markers) motion, respectively obtained by solving the effective master equation (1) and the full master equation (7) including spin and motion. We consider $\eta=0.02$ and $\delta\nu=10^{-3}\gamma_{0}$.
@@ E.3 Coupling to other directions of motion
We consider the same setup as in the main text: The atoms are arranged along the $x$-axis and driven by a classical plane-wave field propagating along $\mathbf{z}$, with drive parameters $\Omega_{j}=\Omega$ and $\Omega_{j}^{(\alpha)}=i\Omega\delta_{\alpha z}$. In this configuration the laser directly cools only the motion along the $z$-axis. The dissipative and coherent couplings between motion along different directions arise from two different processes. The first is recoil heating, and vanishes since $G_{ij}^{\alpha\beta}=0$ for all pairs of directions $\alpha\neq\beta$ [see Eq. (C.2)]. The second is the interference between laser- and dipole-induced spin-motion coupling to first-order in $\eta$, given by Eq. (C.1). Notably, $\Omega_{j}^{(\alpha)}$ and $G_{ij}^{(\beta)}$ are only nonzero for $\alpha=z$ and $\beta=x$, respectively. Thus, motion along $\mathbf{z}$ only couples to motion along $\mathbf{x}$ (through terms of the form $\Omega_{j}^{(z)}G_{ij}^{(x)}$), while motion along $\mathbf{y}$ remains decoupled. For trap frequencies $\nu\gg\gamma_{0}$, this residual coupling is much smaller than the cooling rates along $\mathbf{z}$, $\mathcal{R}_{iziz}^{-}$, and thus it does not impact the cooling dynamics. When $\nu\lesssim\gamma_{0}$, such couplings can still be suppressed under the rotating wave approximation by applying slightly different trap frequencies along $\mathbf{x}$ and $\mathbf{z}$. In this regime, motional degrees of freedom along different directions evolve independently, and enhanced cooling along $\mathbf{z}$ remains possible under the array cooling conditions discussed in the main text. On the other hand, motion along $\mathbf{x}$ and $\mathbf{y}$ can still experience heating, either from recoil (for both directions) or from dipole-induced spin-motion coupling (for $\mathbf{x}$). However, since these heating rates are typically much smaller than the cooling rate along $\mathbf{z}$, they do not lead to appreciable heating during the cooling process.
Figure 11: Two atoms separated by $d=0.2\lambda_{0}$ along the $x$ axis, polarized along the $y$-axis and subject to a three-dimensional harmonic confinement with trapping frequencies $\nu_{1x}=\nu_{1y}=\nu_{1z}=20\gamma_{0}$ for the first atom and $\nu_{1x}=\nu_{1y}=\nu_{1z}=20.01\gamma_{0}$ for the second atom. A drive with Rabi frequency $\Omega=0.1\gamma_{0}$ is applied along the $z$ direction, and we further consider $\eta=0.02$. We plot the average photon number over time along all three directions of motion, $\bar{n}_{\mathrm{\alpha,3D}}$ with $\alpha\in\{x,y,z\}$. The phonon number along the $z$ direction, $\bar{n}_{\mathrm{z,3D}}$, is identical to the case where motion along the $x$ and $y$ directions is neglected, $\bar{n}_{\mathrm{z,1D}}$.
@@ Appendix F Numerical optimization of the cooling rate
Figure 12: (a) Steady-state phonon number $\bar{n}$ and (b) cooling rate $\Gamma$ as a function of drive strength $\Omega$ for two atoms separated by a distance $d=0.2\lambda_{0}$ along the $x$ axis, polarized along the $y$ axis and moving along the $z$ axis. The inset shows an instance of the exponentially decaying phonon number over time for $\Omega=0.2\gamma_{0}$, from which the cooling rate is fitted. (c) Steady state phonon number versus cooling rate for various spacings. Each point corresponds to a different value of $\Omega$. The square marker corresponds to the critical cooling rate $\Gamma_{c}$, obtained at the drive strength that results in a steady state phonon number $10\%$ larger than the optimal one. The triangular marker corresponds to the alternative critical cooling rate $\Gamma_{c}^{\prime}$, obtained at the drive that results in a largest cooling rate while mantaining a steady state phonon number smaller than $1.1\bar{n}_{\textit{ind}}$. (d-e) Critical cooling rates $\Gamma_{c}$ and $\Gamma_{c}^{\prime}$ and their corresponding steady state phonon numbers for chains of $N$ atoms and various spacings. In all panels, we consider $\eta=0.02$ and $\nu_{j}=\bar{\nu}+\delta\nu(j-\lfloor N/2\rfloor-1)$ with $\bar{\nu}=20\gamma_{0}$ and $\delta\nu=5\times 10^{-3}\gamma_{0}$.
@@
Let us consider as an example two atoms in the array cooling regime as defined in Eq. (32) and where the subradiant cooling sideband are resolved (i. e., $\gamma_{A}\ll\bar{\nu}$). The steady state phonon number $\bar{n}$ and the cooling rate $\Gamma$ at which it is reached depend on the drive strength. For sufficiently small drive strength, such that the system remains in the array cooling regime $\Omega^{2}\ll\delta\nu\gamma_{A}/2\eta^{2}$ and the spin-motion coupling is weak $\Omega\ll\gamma_{A}/\eta$, the phonon number exponentially decays to its steady state value at a rate $\Gamma=\mathcal{R}^{-}-\mathcal{R}^{+}\approx\mathcal{R}^{-}\approx 2\eta^{2}\Omega^{2}/\gamma_{A}$, as shown in Fig. 12(a) and (b) for $\bar{\nu}=20\gamma_{0}$. As the drive strength increases, either the array cooling condition or the weak spin-motion coupling condition breaks down, and the effective master equation resulting in Eq. (16) stops being valid. In that case, $\bar{n}$ and $\Gamma$ can only be obtained by numerically solving the full system including spin and motional degrees of freedom. More precisely, $\Gamma$ is obtained by fitting an exponential decay to the phonon number over time, as shown in the inset of Fig. 12(a). We find that, as soon as one of the two conditions is violated, the phonon number starts to increase with increasing $\Omega$ and the cooling rate eventually decreases. To better understand this process, we plot $\bar{n}$ as a function of cooling rate in Fig. 12(c). That is, each data point in this plot corresponds to a different value of $\Omega$. In the main text, we defined the critical cooling rate $\Gamma_{c}$ as the one obtained at the driving strength $\Omega$ that leads to a steady state phonon number $10\%$ larger than the optimal one. For two atoms at a distance $d=0.1\lambda_{0}$, this corresponds to the squared marker in Fig. 12(c). However, this condition is more restrictive for lattices with darker spin modes than attain lower phonon numbers. In other words, we could drive the two atoms at a distance $d=0.1\lambda_{0}$ stronger and thereby attain a larger cooling rate while still reaching steady state phonon numbers smaller than in the case of independent atoms, $\bar{n}<\bar{n}_{\textit{ind}}$. This motivates us to introduce an alternative definition for the critical cooling rate $\Gamma_{c}^{\prime}$ as the largest possible rate that can be reached such that the phonon number remains smaller than $1.1\bar{n}_{\textit{ind}}$. As an example, $\Gamma_{c}^{\prime}$ and its corresponding $\bar{n}^{\prime}$ are shown by the triangular marker in Fig. 12(c) for $d=0.1\lambda_{0}$. In Figs. 12(d) and (e), we plot $\bar{n}$ and $\Gamma_{c}$ for chains of $N$ atoms with various spacing $d$ for $\bar{\nu}=20\gamma_{0}$. Again, squared and triangular markers denote the two different definitions for the critical cooling rate. In all cases, $\Gamma_{c}^{\prime}\gtrsim 2\Gamma_{c}$, showing that the critical cooling rates presented in the main text can be sped up by driving the system stronger at the expense of reaching larger steady state phonon numbers $\bar{n}^{\prime}$. This suggests the following protocol for efficient cooling: first, the system is cooled at $\Gamma^{\prime}$ to a phonon number close to $\bar{n}^{\prime}$; then, the drive is dynamically reduced until the lowest possible phonon number $\bar{n}$ is eventually reached.
@@ Appendix G Cooling $N$ atoms in different geometries
Figure 13: Steady-state phononon number $\bar{n}$ for arrays of different geometries as a function of atom number NN, for (a) $\bar{\nu}=20\gamma_{0}$ and (b) $\bar{\nu}=0.25\gamma_{0}$. We consider a linear ordered chain, a square two-dimensional lattice and a ring of atoms. The atomic chain is placed along the $x$ axis and the atoms are polarized along the $y$ direction. In all cases, the atoms move along the $z$-direction. The squared array and the ring are placed in the xyxy plane and the atoms are circularly polarized in the xyxy plane. For the atomic chain and the squared array, we consider a lattice spacing $d=0.2\lambda_{0}$, while the ring has a radius such that neighboring atoms are at a distance $d=0.2\lambda_{0}$. The results are obtained for $\eta=0.02$ and all atoms have different trap frequencies in increments of $\delta\nu=5\times 10^{-4}\gamma_{0}$, such that the system is in the array cooling regime.